OpenStudy (okdutchman7):

Please help. WILL MEDAL! cos3θ−1=0 @mathstudent55 @Kainui @sleepyjess @satellite73 @phi @zepdrix @robtobey

OpenStudy (okdutchman7):

\[\cos 3\theta-1=0\]

OpenStudy (okdutchman7):

\[\cos3\theta=1\]

OpenStudy (okdutchman7):

This is where I get stuck.

zepdrix (zepdrix):

sec brb :)

zepdrix (zepdrix):

If it was \(\large\rm \cos x=1\), would you know how to proceed?

OpenStudy (okdutchman7):

I think so.

OpenStudy (okdutchman7):

It would be 0,

zepdrix (zepdrix):

Good, cosine is the full length, 1 at an angle of 0. So we would say x=0.

OpenStudy (okdutchman7):

\[0 \le \theta < 360\]

OpenStudy (okdutchman7):

I have to solve for all of these^ values.

zepdrix (zepdrix):

So when \(\large\rm \cos(\color{orangered}{x})=1\), we determined that \(\large\rm \color{orangered}{x}=0\) Let's be a little more specific though, we could add any number of full rotations to this angle 0, and still land at the same spot,\[\large\rm \color{orangered}{x}=0+2k \pi\]Ok with that weird step?

OpenStudy (okdutchman7):

I don't understand.

zepdrix (zepdrix):

cos(0) = 1 cos(0+2pi) = 1 cos(0+4pi) = 1 ... cos(0+2kpi) = 1 The k is some integer, it's some amount of 2pi's.

OpenStudy (okdutchman7):

Okay I got it.

zepdrix (zepdrix):

you determined that 0 angle gives us this 1 length.